A stock that currently trades for $40 per share is expected to pay a year-end dividend of $2 per share. The dividend is expected to grow at a constant rate over time. The stock has a beta of 1.2, the risk-free rate is 5 percent, and the market risk premium is 5 percent. What is the stock’s expected price seven years from today? *
1
Expert's answer
2020-05-11T12:55:00-0400


1.find the rate of return

"RE = Rf + β\times(Rm - Rf)=5+1.2\times5=11"


2.find stock price after 7 years

"FV=PV\times(1+r)^n=40(1+0.11)^7=83.05"